A Complete Bibliography of Publications in Isis, 1910–1919

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 0.03

Title word cross-reference

1496 [32]. 1679 [119]. 1800 [11]. 1884 [7]. 1911 [16]. 1912 [17]. 1er [17]. 1st [17].

Aborigenes [78]. according [155]. Action [39, 72]. actuelle [87]. actuelles [107, 84]. Acvaghosa [8]. Aegypten [30]. Aether [126]. Agnes [42]. al [11]. Albert [45]. alcune [51]. Algerien [30]. alla [51]. allo [28]. Alphonse [15]. Altertum [94]. americaine [75]. Amici [10]. among [78]. analyse [96]. analytique [101, 49, 81, 17]. Ananda [153, 154]. Anaxagora [55]. ancient [55, 56]. année [16]. Annotated [17]. Annuaire [47]. ans [74]. antike [56]. antiquite [62, 7]. Antiquities [156]. Anton [64]. Antonio [10]. août [167]. Arabic [52]. Arabische [30]. arabischen [52]. Arber [42, 42]. Ardigo [70]. Aristoteles [29, 112]. Aristotle [61]. Armin [100, 136, 166]. art [107, 87]. Arts [154]. Ascanio [53]. Astronomical [161]. ateliers [92]. atomism [56]. Atomistik [56]. augmenter [25, 57]. August [112, 69]. Auguste [162, 37]. Australia [78]. Avant [144]. Avant-propos [144]. Awakening [64].

B [121, 39, 72]. B. [82, 119]. B.-L. [82]. Back [143, 169]. Badarayana [155]. Balch [160]. Ball [150, 158]. Barnett [64, 156]. Bateson [129]. Beasts [115]. Bees [115]. Beg [158]. begrundet [166]. Belief [78]. Bensaude [91]. Bernardino [117]. Berthelot [163]. Beuchat [75]. Bhartrihari [63, 63]. Bianchi [121]. Bibliographie [101, 142, 49, 81, 17, 109, 167, 16]. Bibliographique [45]. bibliography [17]. Bignami [121]. Bilancioni [121]. biologischen [127]. Birds [115]. Biringuccio [106]. Birkner [134]. Birma [151]. birth [53]. Bloch [44]. Blutenland [9]. Book [61, 118, 159, 94, 40, 41, 46, 80, 162, 131, 132, 75, 98, 120, 121, 123, 128, 35, 39, 119, 72, 42, 126, 91, 8, 9, 28, 60, 64, 65, 70, 79, 90, 97, 117, 141, 165, 29, 63, 122, 157, 115, 30, 112, 12, 77, 78, 92, 93, 99, 100, 13, 15, 16, 31, 32, 33, 34, 36, 37, 38, 43, 44, 45, 47, 48, 59, 62, 66, 67, 68, 69, 71, 73, 74, 76, 113, 114, 116, 124, 125, 127, 129, 130, 133, 134, 135, 136, 137, 138, 139, 140, 151, 152, 153, 154, 155, 156, 160, 166, 14, 150, 158, 161, 164, 95, 96, 163]. Book [48]. booke [118]. Botanical [54]. Botany [42, 43]. Boutroux [96]. Brabant [31]. Brahma [155]. Brahma-Sutras [155]. Brahmanism [155]. Britain [86]. Britannica [48]. British [43]. Bronnle [64]. Brosse [54]. Brown [40]. Brownian [3]. browniano [3]. Brunschvicg [95]. Buch [9]. Buddhismo [8]. Buffon [120, 120]. Burma [152]. Bussey [35]. Byng [64].

Caillet [45]. Campbell [40]. Candolle [15]. Cankara [155, 155]. Cantideva [64]. Carl [46]. Carlo [8]. Carman [35]. Carmen [22]. Carteggio [121]. Cartesianern [85]. Cartesians [85]. Cartésien [105]. Catalogue [150, 158]. Catalogues [18]. Catherine [118, 14]. Cauchy [88]. Cecil [14]. centenario [53]. centenary [53]. Centuries [67]. Century [164, 54]. Ceylon [151, 154, 152]. Chamberlain [68]. Chapter [42]. Charles [34, 155]. Chemical [146, 85]. Chemie [56]. chemischen [85]. Chemistry [40, 56, 55]. chez [108]. chimica [11, 55]. Chinas [9]. Chisholm [48]. Christ [31]. Christiaan [33]. Christian [150]. Chronique [26, 111, 149, 58, 89, 6]. Chu [64]. cinquante [74]. civilisation [59]. Civilization [145]. Claparede [93]. Clements [159]. Cohen [38]. cokerye [118]. Colding [69]. collections [109, 110]. Collins [64]. Colloquies [159]. Comment [25, 57]. Commentary [155]. Committee [124]. Compendium [155]. Comte [162, 37]. con [51, 121]. Conception [130]. Conceptions [88]. concernenti [120]. Conde [159]. conquète [59]. Continuity [88]. Coomaraswamy [153, 154]. Copernic [163, 113]. correspondance [26, 111, 149, 58, 89, 6]. corrispondenti [10]. cosmologiques [113]. Crafts [154]. Cranmer [64]. Cranmer-Byng [64]. creationnisme [128]. critique [142, 167]. Cruce [160]. current [107, 87, 84]. Cynée [160].

D [64, 156]. Dahlmann [76]. Dahlmann-Waitz [76]. dal [11]. d’Alexandrie [59]. Dampier [14]. Dannemann [125]. dans [7, 107, 77, 62]. Danske [69]. d’après [82]. d’archeologie [75]. dargestellt [123]. David [157]. Day [40]. Dead [78]. découvertes [91]. Definite [88]. del [3, 8]. della [51, 120, 53, 86, 28, 97]. Denis [82]. Denkokonomie [80]. Deonna [99]. depuis [17]. Descartes [85, 85, 108]. Determinants [104]. Deussen [155]. deutschen [46, 76]. Deutsches [36]. deux [110, 74]. development [55]. Diderot [82]. Direction [92]. disciple [71]. Discovery [67]. d’Isis [20]. d’occasion [18]. doctrines [113]. Dogmatics [155]. Dorveaux [66]. d’ouvrages [18]. Drugs [159]. Dschuang [9]. Dsi [9, 9]. Duhem [113]. d’une [16]. Durning [14]. Dyck [36].

E. [73]. Earliest [40]. East [61, 64]. écrivains [163]. Ed [93, 106]. Edward [67, 150, 158]. Edwin [64]. égyptienne [59]. égyptiens [60]. Eighteenth [67]. Eiimcke [114]. Electrical [124]. Electricity [126]. elettromagnetiche [165]. Em. [127]. Emeric [160]. Emil [132]. Energieprinzip [80]. English [24]. engraved [82]. Enquête [93]. Enriques [79]. Enseignement [93]. Entwicklung [125]. époque [91]. Ernst [38]. Erstlinge [32]. esatte [90]. Essai [16]. Etapes [95]. Ethnologie [16]. Eugene [157, 74]. Evolution [42]. exactes [7, 62]. experimental [106]. Exsteens [98].

F [123, 122, 78, 92, 100, 43, 114, 136, 166, 150]. F. [115]. F.-W [92]. Fabre [71]. Favaro [10]. Federigo [79]. Fehr [93]. fenomeni [165]. Ferckel [31]. Ferdinand [134]. Fermat [34]. Ficalho [159]. filosofia [28]. Fisica [97]. Flournoy [93]. forci [69]. Forke [64]. Formichi [8]. Forth [155]. Frances [118]. Francesco [12]. Francis [164]. Frazer [78]. French [82, 168, 107, 87, 17, 2, 84, 148, 108]. Frere [118]. Friedrich [125]. Frischeisen [139]. Frischeisen-Kohler [139]. Front [1, 19, 50, 83, 103, 170]. Frontispice [82]. Frontispiece [82]. Function [88].

G [94, 121, 65, 30, 78, 71, 161]. G. [135]. Gabius [80]. Galen [114]. Galilei [10]. Galileo [10]. Garcia [159]. Garden [64]. Gaston [163]. Gelehrten [15]. Generazione [120]. Genetics [129]. Genie [23]. Genius [23]. Geographical [67]. Geological [148]. Géologique [148]. geometric [108]. geometrique [108]. Geometry [21]. Georg [36]. Georgii [116]. Gerland [73]. German [56, 85, 23, 5, 52]. Gertrude [35]. Geschichte [56, 15, 76, 46, 73, 127]. Geschichts [135]. Geschichts- [135]. Gesprache [9]. Giles [64]. Gino [90]. Giovanni [121, 141]. Gjerulff [69]. glorie [86]. glory [86]. Goethe [122, 68]. Granbretagna [86]. grandes [91]. grands [163]. Graphische [32]. gravé [82]. Great [86]. Grecia [90]. Grosse [13]. Guareschi [12, 11]. Guarino [22]. guérison [138]. Guinea [78]. Guitard [74]. Gustave [59]. Guy [54]. Gynaekologie [31].

H [94, 75, 122, 71, 114, 150]. H.-G [94]. Halbinsel [151]. Handbook [152]. Hans [100, 136, 166]. Heawood [67]. Helmholt [100, 136, 166]. Henri [93]. Henricus [38]. Henriquez [82]. Henry [34]. Herbals [42]. Hindus [21, 146]. Hirschfeld [137]. Histoire [101, 142, 113, 2, 110, 167, 18, 17, 84, 77, 59]. Historisch [122].

Historisch-kritische [122]. History [40, 42, 126, 157, 67, 56, 3, 17, 2, 84]. Hoernes [131]. Hoff [38]. Homosexualitat [137]. Hope [164]. Houston [68]. Hubert [37]. Hugh [48]. Hugo [133, 41]. Human [14]. humanite [25, 57]. Huygens [33]. hypertranscendance [108].

Ibn [64]. Icilio [11, 12]. ihrem [125]. ihrer [125]. II [62]. IIIe [81]. im [94, 114]. Immortality [78]. India [152, 154, 159, 156]. Indian [147]. Indiana [28]. Indien. [151]. inedito [121]. Information [102]. inheritance [23]. Integral [88]. intellectuel [25, 57]. Internationale [47]. interscendance [108]. Introduzione [28]. Inventione [51]. Ioachimi [116]. Ioannis [116]. Islands [78]. Israel [64]. Italia [11]. Italian [22, 51, 53, 3, 86, 55, 106]. IV [101]. Iwan [44].

J [128, 63, 122, 69, 71]. J.-H [71]. Jacobus [38]. Jacques [130]. Jahrbucher [139]. Jahren [32]. Jai [161]. James [40]. January [17]. janvier [17]. Japanese [157, 104]. Java [151]. Jemen [30]. Jequier [59]. Joaquim [91]. Johnston [155]. Joseph [116]. Jourdain [119, 39, 72]. Julien [35]. Julius [29]. jusqu’à [16]. jusqu’en [167].

K. [153, 154]. Kaiser [132]. Kapadia [64]. Karl [32]. Kauffmann [41]. Kaye [65, 161]. Kelka [69]. Kelvin [124]. Kennedy [63]. King [9]. Knobel [150, 158]. Knowledge [146]. Kohlbrugge [122]. Kohler [139]. Kongelige [69]. kritisch [123]. kritische [122]. Kuhner [123]. Kultur [131]. Kungfutse [9].

L [64, 45]. L. [82, 128]. Lamarck [123]. Lanessan [128]. Lao [9]. L’Archeologie. [99]. L’astronomie [91]. Laurent [163]. Least [39, 72]. Leben [123, 38, 114]. Lebens [5]. Lebensbeschreibungen [36]. Legros [71]. Lehre [123]. Leibniz [108]. leibnizienne [108]. Leon [95]. leur [77]. Lia [9]. libri [116]. Lieh [64]. life [5, 130]. Light [64, 33]. L’interscendance [108]. Lionel [64]. Lionell [156]. Literatur [52]. literature [52]. livre [66]. livres [109]. Loeb [130]. Lones [61]. Loo [82]. Lord [124]. Loria [90]. Louis [82]. Ludvig [69]. Luigi [28].

M [63, 96]. Machine [35]. magnetici [165]. Magnus [137]. Magrini [165]. Makers [43]. malayische [151]. Man [35]. Mann [162]. Manner [13]. Mannes [137]. Manuel [75, 45]. Margaret [118]. Markham [159]. Marvin [164]. Masaryk [135]. matematiche [86]. mathematical [86]. mathématiciens [93]. Mathematics [147, 84, 157]. Mathematik [94]. mathématique [93, 95, 96]. mathématiques [84]. Matières [168]. Matter [1, 19, 50, 83, 143, 103, 169, 170, 168]. Matthew [118]. Maurice [98, 77]. Max [139]. Mécanisme [105]. Mechanics [119]. Mechanistic [130]. médecines [66]. Melanesia [78]. Memoires [7, 62]. Men [24]. Mensch [133]. Menschheit [134]. meteoroscopiis [116]. method [106]. méthode [93]. méthodes [99]. Méthodique [168]. metodo [106]. Mettrie [35]. Meyer [114]. Meyer-Steineg [114]. Michel [82]. microscopiche [120]. Mikami [157]. Mind [14]. Mineralogie [52]. Mineralogy [52]. Mittelalter [94]. modern [55]. moderna [55]. Monde [113]. Moret [60]. Morgagni [121]. Moritz [131]. motion [3]. movimento [3]. Mrs [42]. Museum [36]. Mystères [60].

N [65]. nascita [53]. natura [117]. Natural [61]. naturaliste [71]. Nature [39]. Naturforscher [122]. Naturphilosophie [46]. Naturwissenschaften [125]. nautique [91]. Needham [120]. nel [53]. nell’antica [90]. nelle [165]. neueren [56]. Neuzeit [127]. newe [118]. Newell [42]. Newton [119]. Niccolo [51]. Nota [3]. Note [4, 3]. notion [108]. nouveau [160].

Obermaier [133]. Observatories [161]. occidentales [77]. Occultes [45]. Oeuvres [34]. Offray [35]. Old [146]. Oliver [43]. opera [12]. opere [51]. Organisation [101, 142, 27, 167]. Origin [88, 42]. origines [107, 87, 4, 59]. origins [107, 87]. Orta [159]. osservazioni [120]. Ostwald [13, 162].

P [80, 64, 33, 69]. P. [138]. Pagine [70]. Papini [141]. Paracelsus [5]. Paracelsus. [5]. Parker [118, 118]. particolare [51]. particular [51]. parues [17]. Path [64]. Paul [34, 62, 66, 155, 7]. Personlichkeit [123]. Peters [150]. Pflanzennamen [30]. Ph [39]. pharmaceutique [74]. pharmacie [74]. Philip [119, 72]. Philosophie [142, 9, 167, 95, 139]. Philosophy [54]. Physik [73]. physiologie [105]. Pierre [96, 113]. Pitoni [97]. Platearius [66]. Platon [113]. Pleasure [64]. Plinius [112]. poeta [8]. politique [77]. ponderibus [22]. portée [98]. Portrait [82]. Portugal [91]. pragmatismo [141]. précurseur [87]. precursor [87]. prehistoire [98]. Present [40]. presse [74]. pri [69]. principes [96]. Principle [39, 72]. Principles [119]. printing [51]. Prizentita [69]. Problems [129]. progrès [77, 140]. proper [118]. propos [144]. Prostitution [44]. Psychiques [45]. Ptolemy [150, 150]. publications [101, 142, 17, 167]. published [17].

quatuor [116]. quellenden [9]. Quellenkunde [76]. Quelques [107].

R [65, 161]. R. [16]. Radl [127]. Ramanujacarya [65]. rapport [77]. Rassen [134]. razionalismo [79]. reference [51]. reflections [107]. reflexions [107]. Reichenbach [36]. related [17]. relations [107]. relatives [101, 142, 17, 110, 167]. Religion [9]. religions [77]. Religionsphilosophie [135]. rendement [25, 57]. Rene [37]. Reports [124]. rerum [117]. Researches [61]. Review [61, 118, 159, 94, 40, 41, 46, 80, 162, 131, 132, 75, 98, 120, 121, 123, 128, 35, 39, 119, 72, 42, 126, 91, 8, 9, 28, 60, 64, 65, 70, 79, 90, 97, 117, 141, 165, 29, 63, 122, 157, 115, 30, 112, 12, 77, 78, 92, 93, 99, 100, 13, 15, 16, 31, 32, 33, 34, 36, 37, 38, 43, 44, 45, 47, 48, 59, 62, 66, 67, 68, 69, 71, 73, 74, 76, 113, 114, 116, 124, 125, 127, 129, 130, 133, 134, 135, 136, 137, 138, 139, 140, 151, 152, 153, 154, 155, 156, 160, 166, 14, 150, 158, 161, 164, 95, 96, 163]. revues [109, 110]. Rhetici [116]. Richard [9]. riguardo [51]. Rinaldo [97]. Roberto [70]. Royds [115]. Ruska [29]. russischen [135].

S [64, 16]. Sa [99]. Saggio [120]. Saintyves [138]. Satakas [63]. Sayings [63]. scelte [70]. Schriften [123]. Schweinfurth [30]. Science [101, 142, 2, 167, 24, 18, 27, 4, 17, 61, 14]. Sciences [7, 45, 62, 110]. scientifica [12]. scientifiques [7, 62, 163]. Scienza [79]. scienze [90]. Sein [38, 162]. Seine [123]. seinen [123]. seines [5]. Selmi [12]. Selskab [69]. Series [64]. service [74]. ses [107, 99]. Set [155]. Seventeenth [67, 54]. Seventeenth-Century [54]. Siam [151]. siècle [105]. siècles [74]. Siegel [46]. Signori [120]. Silvanus [33]. Silvio [165]. Simples [159, 66]. since [17]. Singh [161]. sistema [120]. sketch [5]. Skizze [5]. Smith [157]. Sobrero [53, 53]. social [77]. sociale [77]. Soixante [110]. Soixante-deux [110]. some [51, 107]. Soul [64]. Spallanzani [120]. Spampanato [117]. sperimentale [106]. sphaericis [116]. Sridharacaya [65]. stampa [51]. Standards [124]. Standpoint [155]. Stars [150, 158]. Steier [112]. Steinbuch [29]. Steineg [114]. Steinmetz [16]. Stewart [68]. storia [3, 97]. Straits [78]. Studien [122]. studio [28]. Study [54]. style [82]. sua [53, 12]. Suali [28]. Subject [168]. Sudhoff [32]. sudlichen [9]. sue [51]. Sul [141]. sulla [3]. Sutras [155]. sviluppo [55]. Sydney [164]. Synthèse [148]. synthesis [148]. synthetique [109]. Syphilisliteratur [32]. System [155]. systématique [16]. Systeme [113].

’t [38, 115]. T. [126]. Table [168]. Tag [114]. Tannery [7, 34, 62]. Tao [9]. Taoist [64]. Tartaglia [51]. Taylor [92]. te [9]. Teachings [64]. Telesio [117]. tendances [107, 84]. teoria [55]. teorie [165]. tezi [69]. Th [93, 135]. Their [42]. Theodor [114]. theorie [107, 87]. Theorien [85, 127]. Theories [126, 85]. Theory [104, 107, 87, 55]. Thereon [155]. Thomas [61, 31, 160]. Thompson [33]. Thought [54]. Three [24]. till [40]. Tille [100, 136, 166]. Times [40]. Tome [62]. Torres [78]. tous [98]. toutes [101, 142, 167]. transcendance [108]. transcendence [108]. Transformisme [128]. Travagliata [51]. travail [93]. Travellers [152]. Treatise [33]. trends [107, 84]. Triangulis [116]. Trisatika [65]. Troilo [70]. Troubled [51]. Tse [9]. Tufail [64]. Typographische [32]. Tzu [64].

Ulugh [158]. un’antica [55]. Union [74]. Urgrund [9]. Urkunden [36]. Urzeit [131]. use [55]. utilizzazione [55].

Valenzlehre [41]. valeur [99]. Validity [39]. Vannoccio [106]. varie [165]. Ve [142]. Vedanta [155]. Vererbung [23]. Verneri [116]. Vernes [77]. Veronese [22]. verrues [138]. Videnskabernes [69]. Vie [47, 71, 167]. Vincenzio [10]. Vincenzo [117]. Virgil [115]. Visvakarma [153]. Viviani [10]. Vol [78]. Volker [134]. Volume [102]. vom [123, 9]. Vorderindien [151]. Vorzeit [132, 133].

W [92, 43]. wahre [9]. Waitz [76]. Waldemar [99]. Walther [36]. War [145]. Weibes [137]. Weltgeschichte [100, 136, 166]. Werk [162]. Wesentliche [123]. Whetham [14]. Whittaker [126]. Wilhelm [162, 9, 13]. William [14, 129]. Willing [160]. Wirken [38]. Wisdom [64]. Wise [63]. Wissenschaften [15]. works [51]. Worship [78]. Wurschmidt [116].

XVIIe [105]. XXIX [10].

Yang [64]. Year [48]. Year-Book [48]. Yoshio [157].

Zeuthen [94]. Zur [15, 135]. Zusammenhange [125].

References

Anonymous:1913:FMa

[1] Anonymous. Front matter. Isis, 1(1):1–2, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

     org/stable/223803.

Sarton:1913:LSF

[2] George Sarton. L’histoire de la science. (French) [The history of science]. Isis, 1(1):3–46, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print),

     1545-6994 (electronic). URL http://www.jstor.org/stable/223804.

Guareschi:1913:NSS

[3] I. Guareschi. Nota sulla storia del movimento browniano. (Italian) [Note on the history of Brownian motion]. Isis, 1(1):47–52, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

     www.jstor.org/stable/223805.

Milhaud:1913:NOS

[4] G. Milhaud. Note sur les origines de la science. Isis, 1(1):53–61, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

     URL http://www.jstor.org/stable/223806.

Radl:1913:PSS

[5] Em. Radl. Paracelsus. Eine Skizze seines Lebens. (German) [Paracelsus. A sketch of his life]. Isis, 1(1):62–94, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223807.

Sarton:1913:CC

[6] George Sarton. Chronique et correspondance. Isis, 1(1):95–113, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223808.

Anonymous:1913:BSS

[7] Anonymous. Memoires scientifiques. I. Sciences exactes dans l’antiquité, 1876–1884 by Tannery, Paul. Isis, 1(1):114–115, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223809.

Masson-Oursel:1913:BRBa

[8] P. Masson-Oursel. Book review: Acvaghosa, poeta del Buddhismo by Carlo Formichi. Isis, 1(1):115–117, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223810.

Masson-Oursel:1913:BRBb

[9] P. Masson-Oursel. Book review: Die Religion und Philosophie Chinas by Richard Wilhelm; Gesprache by Kungfutse; Tao te King by Lao Tse; Das wahre Buch vom quellenden Urgrund by Lia Dsi; Das wahre Buch vom sudlichen Blutenland by Dschuang Dsi. Isis, 1(1):117–120, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223811.

Anonymous:1913:ACG

[10] Anonymous. “Amici e corrispondenti di Galileo Galilei XXIX, Vincenzio Viviani” by Favaro, Antonio. Isis, 1(1):120–122, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223812.

Anonymous:1913:BCI

[11] Anonymous. “La chimica in Italia dal 1750 al 1800” by Guareschi, Icilio; I. Guareschi. Isis, 1(1):122, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223813.

S:1913:BRBa

[12] G. S. Book review: “Francesco Selmi e la sua opera scientifica” by Guareschi, Icilio. Isis, 1(1):123–124, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223814.

S:1913:BRBb

[13] G. S. Book review: “Grosse Manner” by Ostwald, Wilhelm. Isis, 1(1): 124–125, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223815.

Sarton:1913:BRB

[14] George Sarton. Book review: Science and the Human Mind by Whetham William Cecil Dampier; Whetham Catherine Durning. Isis, 1(1):125– 132, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223816.

S:1913:BRBc

[15] G. S. Book review: Zur Geschichte der Wissenschaften und der Gelehrten by Candolle Alphonse de. Isis, 1(1):132–133, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223817.

S:1913:BRBd

[16] G. S. Book review: Essai d’une bibliographie systématique de l’Ethnologie jusqu’à l’année 1911 by Steinmetz, S. R. Isis, 1(1):133–135, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223818.

Sarton:1913:BAP

[17] George Sarton. Bibliographie analytique des publications relatives à l’histoire, de la science parues depuis le 1er janvier 1912. (French) [Annotated bibliography of publications related to the history of science published since the 1st January 1912]. Isis, 1(1):136–188, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223819.

Anonymous:1913:COO

[18] Anonymous. Catalogues d’ouvrages d’occasion sur l’histoire de la science. Isis, 1(1):189–190, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223820.

Anonymous:1913:FMb

[19] Anonymous. Front matter. Isis, 1(2):191–192, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/223936.

Sarton:1913:B

[20] George Sarton. Le but d’Isis. Isis, 1(2):193–196, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223937.

Smith:1913:GH

[21] David Eugene Smith. The geometry of the hindus. Isis, 1(2):197–204, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223938.

Favaro:1913:CPG

[22] A. Favaro. Il “Carmen de ponderibus,, di Guarino Veronese. (Italian) [the “Carmen de ponderibus” of Guarino Veronese]. Isis, 1(2):205–207, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223939.

Ostwald:1913:GVG

[23] Wilhelm Ostwald. Genie und Vererbung. (German) [genius and inheritance]. Isis, 1(2):208–214, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223940.

Whetham:1913:TEM

[24] C. D. Whetham and W. C. D. Whetham. Three English men of science. Isis, 1(2):215–218, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223941.

Sarton:1913:CARa

[25] George Sarton. Comment augmenter le rendement intellectuel de l’humanite? Isis, 1(2):219–242, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223942.

Anonymous:1913:CC

[26] Anonymous. Chronique et correspondance. Isis, 1(2):243–251, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223943.

M:1913:OS

[27] P. M. Organisation de la science. Isis, 1(2):252–263, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223944.

Masson-Oursel:1913:BRBc

[28] P. Masson-Oursel. Book review: Introduzione allo studio della filosofia Indiana by Luigi Suali. Isis, 1(2):264–266, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223945.

Oursel:1913:BRBa

[29] P. Masson Oursel. Book review: Das Steinbuch des Aristoteles by Julius Ruska. Isis, 1(2):266–267, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223946.

Ruska:1913:BRB

[30] Julius Ruska. Book review: Arabische Pflanzennamen aus Aegypten, Algerien und Jemen by G. Schweinfurth. Isis, 1(2):268–271, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223947.

S:1913:BRBe

[31] G. S. Book review: Die Gynaekologie des Thomas von Brabant by Christ Ferckel. Isis, 1(2):271–272, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223948.

S:1913:BRBf

[32] G. S. Book review: Graphische und Typographische Erstlinge der Syphilisliteratur aus den Jahren 1495 und 1496 by Karl Sudhoff. Isis, 1(2):272– 273, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223949.

S:1913:BRBg

[33] G. S. Book review: Treatise on Light by Christiaan Huygens; Silvanus P. Thompson. Isis, 1(2):273–274, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223950.

S:1913:BRBh

[34] G. S. Book review: Oeuvres de Fermat by Paul Tannery; Charles Henry. Isis, 1(2):274, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/223951.

J:1913:BRBa

[35] J. Book review: Man a Machine by Julien Offray de La Mettrie; Gertrude Carman Bussey. Isis, 1(2):274–275, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223952.

S:1913:BRBi

[36] G. S. Book review: Georg von Reichenbach. Deutsches Museum Lebensbeschreibungen und Urkunden by Walther von Dyck. Isis, 1(2):275–276, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/223953.

S:1913:BRBj

[37] G. S. Book review: Auguste Comte by Rene Hubert. Isis, 1(2):276, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223954.

S:1913:BRBk

[38] G. S. Book review: Jacobus Henricus van ’t Hoff. Sein Leben und Wirken by Ernst Cohen. Isis, 1(2):276–278, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223955.

J:1913:BRBb

[39] J. Book review: The Nature and Validity of the Principle of Least Action by Ph. E. B. Jourdain. Isis, 1(2):278–279, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223956.

Bloch:1913:BRBa

[40] Ernst Bloch. Book review: A History of Chemistry from the Earliest Times till the Present Day by Campbell Brown James. Isis, 1(2):279– 280, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223957.

Bloch:1913:BRBb

[41] Ernst Bloch. Book review: Die Valenzlehre by Kauffmann Hugo. Isis, 1(2): 280–281, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223958.

K:1913:BRB

[42] K and G. S. Book review: Herbals, Their Origin and Evolution. A Chapter in the History of Botany (1470–1670) by Agnes Arber; Mrs. E. A. Newell Arber. Isis, 1(2):281–282, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223959.

S:1913:BRBl

[43] G. S. Book review: Makers of British Botany by F. W. Oliver. Isis, 1(2): 282–284, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223960.

S:1913:BRBm

[44] G. S. Book review: Die Prostitution by Iwan Bloch. Isis, 1(2):284– 285, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223961.

S:1913:BRBn

[45] G. S. Book review: Manuel Bibliographique des Sciences Psychiques ou Occultes by Albert L. Caillet. Isis, 1(2):285–287, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223962.

Bloch:1913:BRBc

[46] Ernst Bloch. Book review: Geschichte der deutschen Naturphilosophie by Siegel Carl. Isis, 1(2):287–289, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223963.

S:1913:BRBo

[47] G. S. Book review: Annuaire de la Vie Internationale. Isis, 1(2):289– 290, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223964.

S:1913:BRBp

[48] G. S. Book review: The Britannica Year-Book by Hugh Chisholm. Isis, 1 (2):290–292, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/223965.

S:1913:BA

[49] G. S. Bibliographie analytique. Isis, 1(2):293–326, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223966.

Anonymous:1913:FMc

[50] Anonymous. Front matter. Isis, 1(3):327–328, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/224133.

Favaro:1913:NTD

[51] Antonio Favaro. Di Niccolo Tartaglia e della stampa di alcune delle sue opere con particolare riguardo alla “Travagliata Inventione”. (Italian) [On Niccolo Tartaglia and the printing of some of his works with particular reference to “Troubled Inventione”]. Isis, 1(3):329–340, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224135.

Ruska:1913:MAL

[52] Julius Ruska. Die Mineralogie in der arabischen Literatur. (German) [Mineralogy in Arabic literature]. Isis, 1(3):341–350, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224136.

Guareschi:1913:ASN

[53] I. Guareschi. Ascanio Sobrero nel centenario della sua nascita. (Italian) [Ascanio Sobrero on the centenary of his birth]. Isis, 1(3):351–358, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224137.

Arber:1913:BPG

[54] Agnes Arber. The botanical philosophy of Guy de la Brosse: A study in Seventeenth-Century thought. Isis, 1(3):359–369, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224138.

Mieli:1913:TAC

[55] Aldo Mieli. La teoria di Anaxagora e la chimica moderna (lo sviluppo e l’utilizzazione di un’antica teoria). (Italian) [The theory of Anaxagora and modern chemistry (the development and use of an ancient theory)]. Isis, 1(3):370–376, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224139.

Bloch:1913:AAN

[56] Ernst Bloch. Die antike Atomistik in der neueren Geschichte der Chemie. (German) [Ancient atomism in the recent history of chemistry]. Isis, 1(3):

377–415, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224140.

Sarton:1913:CARb

[57] George Sarton. Comment augmenter le rendement intellectuel de l’humanite? Isis, 1(3):416–473, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224141.

Mieli:1913:CCa

[58] Aldo Mieli and Ernst Bloch. Chronique et correspondance. Isis, 1(3): 474–502, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224142.

S:1913:BRBq

[59] G. S. Book review: Histoire de la civilisation égyptienne, des origines à la conquète d’Alexandrie by Gustave Jequier. Isis, 1(3):503–504, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224143.

Masson-Oursel:1913:BRBd

[60] P. Masson-Oursel. Book review: Mystères égyptiens by A. Moret. Isis, 1 (3):504–505, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/224144.

Arber:1913:BRB

[61] Agnes Arber. Book review: Aristotle’s Researches in Natural Science by Thomas East Lones. Isis, 1(3):505–509, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224145.

S:1913:BRBr

[62] G. S. Book review: Memoires scientifiques, Tome II. Sciences exactes dans l’antiquite (1883–1898) by Paul Tannery. Isis, 1(3):509–512, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224146.

Oursel:1913:BRBb

[63] P. Masson Oursel. Book review: The Satakas or Wise Sayings of Bhartrihari by J. M. Kennedy; Bhartrihari. Isis, 1(3):512–513, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224147.

Masson-Oursel:1913:BRBe

[64] P. Masson-Oursel. Book review: The Wisdom of the East Series by L. Cranmer-Byng; S. A. Kapadia; The Wisdom of Israel by Edwin Collins; The Awakening of the Soul by Ibn Tufail; P. Bronnle; The Path of Light by Cantideva; L. D. Barnett; Taoist Teachings by Lieh Tzu; Lionel Giles; Yang Chu’s Garden of Pleasure by Anton Forke. Isis, 1(3):513–516, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224148.

Masson-Oursel:1913:BRBf

[65] P. Masson-Oursel. Book review: The Trisatika of Sridharacaya by N. Ramanujacarya; G. R. Kaye. Isis, 1(3):516–517, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224149.

S:1913:BRBs

[66] G. S. Book review: Le livre des simples médecines by Paul Dorveaux; Platearius. Isis, 1(3):517–518, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224150.

S:1913:BRBt

[67] G. S. Book review: A History of Geographical Discovery in the Seventeenth and Eighteenth Centuries by Edward Heawood. Isis, 1(3):518–519, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224151.

S:1913:BRBu

[68] G. S. Book review: Goethe by Houston Stewart Chamberlain. Isis, 1(3): 519–521, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224152.

S:1913:BRBv

[69] G. S. Book review: Kelka tezi pri la forci. Prizentita a det Kongelige Danske Videnskabernes Selskab by Ludvig August Colding; J. P. Gjerulff. Isis, 1(3):522–524, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224153.

Mieli:1913:BRBa

[70] Aldo Mieli. Book review: Pagine scelte by Ardigo Roberto; E. Troilo. Isis, 1(3):524–526, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224154.

S:1913:BRBw

[71] G. S. Book review: La vie de J.-H. Fabre, naturaliste, par un disciple by G. Legros. Isis, 1(3):526–527, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224155.

Jourdain:1913:BRB

[72] Jourdain. Book review: The Principle of Least Action by Philip E. B. Jourdain. Isis, 1(3):527, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224156.

S:1913:BRBx

[73] G. S. Book review: Geschichte der Physik by Gerland E. Isis, 1(3):527– 529, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224157.

S:1913:BRBy

[74] G. S. Book review: Deux siècles de presse au service de la pharmacie et cinquante ans de “Union pharmaceutique” by Eugene Guitard. Isis, 1(3): 529–530, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224158.

Engerrand:1913:BRBa

[75] Jorge Engerrand. Book review: Manuel d’archeologie americaine by H. Beuchat. Isis, 1(3):530–537, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224159.

S:1913:BRBz

[76] G. S. Book review: Quellenkunde der deutschen Geschichte by Dahlmann-Waitz. Isis, 1(3):537–538, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224160.

S:1913:BRBaa

[77] G. S. Book review: Histoire sociale des religions. — I. Les Religions occidentales dans leur rapport avec le progrès politique et social by Maurice Vernes. Isis, 1(3):538–539, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224161.

S:1913:BRBab

[78] G. S. Book review: The Belief in Immortality and the Worship of the Dead. — Vol. I. The Belief among the Aborigenes of Australia, the Torres Straits Islands, New Guinea and Melanesia by F. G. Frazer. Isis, 1(3): 540, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224162.

Mieli:1913:BRBb

[79] Aldo Mieli. Book review: Scienza e razionalismo by Enriques Federigo. Isis, 1(3):541–542, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224163.

Bloch:1913:BRBd

[80] Ernst Bloch. Book review: Denkokonomie und Energieprinzip by P. Gabius. Isis, 1(3):542, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224164.

S:1913:IBA

[81] G. S. Iiie bibliographie analytique. Isis, 1(3):543–574, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224165.

Anonymous:1913:FPD

[82] Anonymous. Frontispice: Portrait de Denis Diderot (1713–1784), gravé par B.-L. Henriquez, d’après Louis Michel van Loo. (French) [Frontispiece: Portrait of Denis Diderot (1713–1784), engraved by B.-L. Henriquez, in the style of Louis Michel van Loo]. Isis, 1(3):??, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224134.

Anonymous:1913:FMd

[83] Anonymous. Front matter. Isis, 1(4):575–576, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/223780.

Sarton:1913:TAH

[84] George Sarton. Les tendances actuelles de l’histoire des mathématiques. (French) [Current trends in the history of mathematics]. Isis, 1(4):577– 589, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223781.

Bloch:1913:CTD

[85] Ernst Bloch and Aldo Mieli. Die chemischen Theorien bei Descartes und den Cartesianern. (German) [The chemical theories of Descartes and the Cartesians]. Isis, 1(4):590–636, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223782.

Loria:1913:GMD

[86] Gino Loria. Le glorie matematiche della Granbretagna. (Italian) [The mathematical glory of Great Britain]. Isis, 1(4):637–654, ???? 1913. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/223783.

Gaffarel:1913:PTA

[87] Jacques Gaffarel and Waldemar Deonna. Un précurseur de la théorie actuelle des origines de l’art. (French) [A precursor of the current theory of the origins of art]. Isis, 1(4):655–660, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223784.

Jourdain:1913:OCC

[88] Philip E. B. Jourdain. The origin of Cauchy’s conceptions of a definite integral and of the continuity of a function. Isis, 1(4):661–703, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223785.

Mieli:1913:CCb

[89] Aldo Mieli. Chronique et correspondance. Isis, 1(4):704–713, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223786.

Mieli:1913:BRBc

[90] Aldo Mieli. Book review: Le scienze esatte nell’antica Grecia by Loria Gino. Isis, 1(4):714–716, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223787.

Mascart:1913:BRB

[91] Jean Mascart. Book review: L’astronomie nautique au Portugal a l’époque des grandes découvertes by Joaquim Bensaude. Isis, 1(4):716–718, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223788.

S:1913:BRBac

[92] G. S. Book review: La Direction des ateliers by F.-W. Taylor. Isis, 1 (4):718, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223789.

S:1913:BRBad

[93] G. S. Book review: Enquête de l’“Enseignement mathématique” sur la méthode de travail des mathématiciens by Henri Fehr; Th. Flournoy; Ed. Claparede. Isis, 1(4):719, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223790.

B:1913:BRB

[94] V. B. Book review: Die Mathematik im Altertum und im Mittelalter by H.-G. Zeuthen. Isis, 1(4):719–721, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223791.

Turriere:1913:BRBa

[95] Emile Turriere. Book review: Les Etapes de la philosophie mathématique by Leon Brunschvicg. Isis, 1(4):721–734, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223792.

Turriere:1913:BRBb

[96] Emile Turriere. Book review: Les principes de l’analyse mathématique by M. Pierre Boutroux. Isis, 1(4):734–742, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223793.

Mieli:1913:BRBd

[97] Aldo Mieli. Book review: Storia della Fisica by Pitoni, Rinaldo. Isis, 1(4): 742–744, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223794.

Engerrand:1913:BRBb

[98] Jorge Engerrand. Book review: La prehistoire à la portée de tous by Maurice Exsteens. Isis, 1(4):744, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223795.

S:1913:BRBae

[99] G. S. Book review: L’Archeologie. Sa valeur, ses méthodes by Waldemar Deonna. Isis, 1(4):745–754, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223796.

S:1913:BRBaf

[100] G. S. Book review: Weltgeschichte by Armin Tille; Hans F. Helmholt. Isis, 1(4):754–756, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223797.

Anonymous:1913:IBA

[101] Anonymous. IV bibliographie analytique de toutes les publications relatives a l’histoire, et a l’organisation de la science. Isis, 1(4):757–791, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223798.

Anonymous:1913:VI

[102] Anonymous. Volume information. Isis, 1(4):792–824, ???? 1913. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223779.

Anonymous:1914:FM

[103] Anonymous. Front matter. Isis, 2(1):5–8, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223552.

Mikami:1914:JTD

[104] Yoshio Mikami. On the Japanese theory of determinants. Isis, 2(1): 9–36, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223553.

Georges-Berthier:1914:MCP

[105] Aug. Georges-Berthier. Le mécanisme cartésien et la physiologie au XVIIe siècle. Isis, 2(1):37–89, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223554.

Mieli:1914:VBE

[106] Aldo Mieli. Vannoccio Biringuccio ed il metodo sperimentale. (Italian) [Vannoccio Biringuccio and the experimental method]. Isis, 2(1):90–99, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/223555.

Deonna:1914:QRT

[107] W. Deonna. Quelques reflexions sur la theorie des origines de l’art dans ses relations avec les tendances actuelles. (French) [Some reflections on the theory of the origins of art in its relations with current trends]. Isis, 2(1): 100–105, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223556.

Turriere:1914:NTG

[108] Emile Turriere. La notion de transcendance geometrique chez Descartes et Leibniz. L’interscendance leibnizienne et l’hypertranscendance. (French) [The geometric notion of transcendence in Descartes and Leibniz. The interscendance of Leibniz and hypertranscendance]. Isis, 2(1):106–124, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223557.

Sarton:1914:BSR

[109] George Sarton. Bibliographie synthetique des revues et des collections de livres. Isis, 2(1):125–131, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223558.

Sarton:1914:SDR

[110] George Sarton. Soixante-deux revues et collections relatives a l’histoire des sciences. Isis, 2(1):132–161, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223559.

E:1914:CC

[111] J. E. and A. Korn. Chronique et correspondance. Isis, 2(1):162–201, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223560.

Russell:1914:BRB

[112] E. S. Russell. Book review: Aristoteles und Plinius by August Steier. Isis, 2(1):202–203, June 1914. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223561.

S:1914:BRBa

[113] G. S. Book review: Le Systeme du Monde. Histoire des doctrines cosmologiques de Platon a Copernic by Duhem, Pierre. Isis, 2(1):203–204, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/223562.

S:1914:BRBb

[114] G. S. Book review: Ein Tag im Leben des Galen by Meyer-Steineg, Theodor; F. H. Eiimcke. Isis, 2(1):204–205, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223563.

Robertson:1914:BRB

[115] D. S. Robertson. Book review: The Beasts, Birds and Bees of Virgil by Royds, T. F. Isis, 2(1):205, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223564.

S:1914:BRBc

[116] G. S. Book review: De Triangulis sphaericis, libri quatuor, De meteoroscopiis by Verneri, Ioannis; Georgii Ioachimi Rhetici; Joseph Wurschmidt. Isis, 2(1):205–206, June 1914. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223565.

Mieli:1914:BRBa

[117] Aldo Mieli. Book review: De rerum natura by Telesio, Bernardino; Vincenzo Spampanato. Isis, 2(1):206–208, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223566.

Arber:1914:BRB

[118] Agnes Arber. Book review: A proper newe booke of cokerye by Frere, Catherine Frances; Matthew Parker; Margaret Parker. Isis, 2(1):208, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223567.

J:1914:BRB

[119] J. Book review: The Principles of Mechanics with Newton, from 1666 to 1679 by Jourdain Philip, E. B. Isis, 2(1):209, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223568.

Georges-Berthier:1914:BRBa

[120] A. Georges-Berthier. Book review: Saggio di osservazioni microscopiche concernenti il sistema della Generazione dei Signori di Needham e Buffon by Spallanzani; Needham; Buffon. Isis, 2(1):209–213, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223569.

Georges-Berthier:1914:BRBb

[121] A. Georges-Berthier. Book review: Carteggio inedito di G. B. Morgagni con Giovanni Bianchi by G. Bilancioni; A. Bignami. Isis, 2(1):213, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223570.

Radl:1914:BRB

[122] Em. Radl. Book review: Historisch-kritische Studien über Goethe als Naturforscher by J. H. F. Kohlbrugge. Isis, 2(1):213–216, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223571.

Georges-Berthier:1914:BRBc

[123] A. Georges-Berthier. Book review: Die Lehre vom Leben. Seine Personlichkeit und das Wesentliche aus seinen Schriften kritisch dargestellt by Lamarck; F. Kuhner. Isis, 2(1):216–217, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223572.

S:1914:BRBd

[124] G. S. Book review: Reports of the Committee on Electrical Standards by Lord Kelvin. Isis, 2(1):217–218, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223573.

S:1914:BRBe

[125] G. S. Book review: Die Naturwissenschaften in ihrer Entwicklung und in ihrem Zusammenhange by Dannemann, Friedrich. Isis, 2(1):218–222, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223574.

Magrini:1914:BRB

[126] Silvio Magrini. Book review: A History of the Theories of Aether and Electricity by Whittaker, E. T. Isis, 2(1):222–224, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223575.

S:1914:BRBf

[127] G. S. Book review: Geschichte der biologischen Theorien in der Neuzeit by Radl, Em. Isis, 2(1):224–229, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223576.

Guinet:1914:BRB

[128] L. Guinet. Book review: Transformisme et creationnisme by De Lanessan, J. L. Isis, 2(1):229–230, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223577.

S:1914:BRBg

[129] G. S. Book review: Problems of Genetics by Bateson, William. Isis, 2(1): 230–231, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223578.

S:1914:BRBh

[130] G. S. Book review: The Mechanistic Conception of Life by Loeb, Jacques. Isis, 2(1):232, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/223579.

E:1914:BRBa

[131] J. E. Book review: Kultur der Urzeit by Hoernes, Moritz. Isis, 2(1): 232–233, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223580.

E:1914:BRBb

[132] J. E. Book review: Aus der Vorzeit by Kaiser, Emil. Isis, 2(1):233– 234, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223581.

S:1914:BRBi

[133] G. S. Book review: Der Mensch der Vorzeit by Hugo Obermaier. Isis, 2 (1):234–237, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/223582.

S:1914:BRBj

[134] G. S. Book review: Die Rassen und Volker der Menschheit by Ferdinand Birkner. Isis, 2(1):237–238, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223583.

S:1914:BRBk

[135] G. S. Book review: Zur russischen Geschichts- und Religionsphilosophie by Masaryk, Th. G. Isis, 2(1):238–239, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223584.

S:1914:BRBl

[136] G. S. Book review: Weltgeschichte by Tille, Armin; Hans F. Helmholt. Isis, 2(1):239–240, June 1914. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223585.

S:1914:BRBm

[137] G. S. Book review: Die Homosexualitat des Mannes und des Weibes by Hirschfeld, Magnus. Isis, 2(1):241–242, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223586.

S:1914:BRBn

[138] G. S. Book review: La guérison des verrues by Saintyves, P. Isis, 2(1): 242–243, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223587.

S:1914:BRBo

[139] G. S. Book review: Jahrbucher der Philosophie by Max Frischeisen-Kohler. Isis, 2(1):243–244, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223588.

S:1914:BRBp

[140] G. S. Book review: Le Progrès. Isis, 2(1):245, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223589.

Mieli:1914:BRBb

[141] Aldo Mieli. Book review: Sul pragmatismo by Papini, Giovanni. Isis, 2(1): 245–247, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223590.

Anonymous:1914:VBC

[142] Anonymous. Ve bibliographie critique de toutes les publications relatives à l’histoire, à la philosophie et à l’organisation de la science. Isis, 2(1): 248–310, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223591.

Anonymous:1914:BM

[143] Anonymous. Back matter. Isis, 2(1):??, June 1914. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223592.

Sarton:1919:AP

[144] George Sarton. Avant-propos. Isis, 2(2):313–314, September 1919. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/223880.

Sarton:1919:WC

[145] George Sarton. War and civilization. Isis, 2(2):315–321, September 1919. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223881.

Ray:1919:CKH

[146] Praphulla Chandra Ray. Chemical knowledge of the Hindus of old. Isis, 2(2):322–325, September 1919. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223882.

Kaye:1919:IM

[147] G. R. Kaye. Indian mathematics. Isis, 2(2):326–356, September 1919. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223883.

Sarton:1919:SGF

[148] George Sarton. La synthèse géologique de 1775 à 1918. (French) [Geological synthesis from 1775 to 1918]. Isis, 2(2):357–394, September 1919. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223884.

F:1919:CC

[149] F. M. F., Aldo Mieli, and G. S. Chronique et correspondance. Isis, 2 (2):395–400, September 1919. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223885.

Sarton:1919:BRBa

[150] G. Sarton. Book review: Ptolemy’s Catalogue of Stars by Ptolemy; Christian H. F. Peters; Edward Ball Knobel. Isis, 2(2):401, September 1919. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223886.

S:1919:BRBa

[151] G. S. Book review: Indien. Ceylon. Vorderindien. Birma. Die malayische Halbinsel. Siam. Java. Isis, 2(2):401–402, September 1919. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223887.

S:1919:BRBb

[152] G. S. Book review: A Handbook for Travellers in India, Burma and Ceylon. Isis, 2(2):402–403, September 1919. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223888.

S:1919:BRBc

[153] G. S. Book review: Visvakarma by Coomaraswamy, Ananda K. Isis, 2 (2):403–404, September 1919. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223889.

S:1919:BRBd

[154] G. S. Book review: The Arts and Crafts of India and Ceylon by Coomaraswamy, Ananda K. Isis, 2(2):404–407, September 1919. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223890.

S:1919:BRBe

[155] G. S. Book review: The System of the Vedanta, according to Badarayana’s Brahma-Sutras and Cankara’s Commentary Thereon Set Forth as a Compendium of the Dogmatics of Brahmanism from the Standpoint of Cankara by Paul Deussen; Charles Johnston. Isis, 2(2):407–408, September 1919. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223891.

S:1919:BRBf

[156] G. S. Book review: Antiquities of India by Lionell D. Barnett. Isis, 2 (2):408–410, September 1919. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223892.

Reymond:1919:BRB

[157] Arnold Reymond. Book review: A History of Japanese Mathematics by David Eugene Smith; Yoshio Mikami. Isis, 2(2):410–413, September 1919. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223893.

Sarton:1919:BRBb

[158] George Sarton. Book review: Catalogue of Stars by Ulugh Beg; Edward Ball Knobel. Isis, 2(2):413–415, September 1919. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223894.

Arber:1919:BRB

[159] Agnes Arber. Book review: Colloquies on the Simples and Drugs of India by Orta, Garcia da; Conde de Ficalho; Clements Markham. Isis, 2(2):415– 418, September 1919. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223895.

S:1919:BRBg

[160] G. S. Book review: Le nouveau Cynée by Cruce, Emeric; Thomas Willing Balch. Isis, 2(2):418–421, September 1919. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223896.

Sarton:1919:BRBc

[161] G. Sarton. Book review: The Astronomical Observatories of Jai Singh by G. R. Kaye. Isis, 2(2):421–423, September 1919. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223897.

Bloch:1919:BRB

[162] Ernst Bloch. Book review: Auguste Comte, der Mann und sein Werk by Wilhelm Ostwald. Isis, 2(2):423–424, September 1919. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223898.

Turriere:1919:BRB

[163] E. Turriere. Book review: Les grands écrivains scientifiques, de Copernic à Berthelot by Gaston Laurent. Isis, 2(2):424–425, September 1919. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223899.

Sarton:1919:BRBd

[164] George Sarton. Book review: The Century of Hope by Francis Sydney Marvin. Isis, 2(2):425–427, September 1919. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223900.

Mieli:1919:BRB

[165] Aldo Mieli. Book review: I fenomeni magnetici nelle varie teorie elettromagnetiche by Silvio Magrini. Isis, 2(2):427–428, September 1919. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/223901.

S:1919:BRBh

[166] G. S. Book review: Weltgeschichte begrundet by Tille, Armin; Hans F. Helmholt. Isis, 2(2):428, September 1919. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223902.

Sarton:1919:VBC

[167] George Sarton. VIe bibliographie critique de toutes les publications relatives à l’histoire, la philosophie et l’organisation de la science (jusqu’en août 1914). Isis, 2(2):429–480, September 1919. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223903.

Anonymous:1919:TMM

[168] Anonymous. Table méthodique des matières. (French) [subject matter table]. Isis, 2(2):481–487, September 1919. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223904.

Anonymous:1919:BM

[169] Anonymous. Back matter. Isis, 2(2):??, September 1919. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223905.

Anonymous:1919:FM

[170] Anonymous. Front matter. Isis, 2(2):??, September 1919. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/223879.